OpenStudy (anonymous):

how to factor 36x^2y^6 - 16

OpenStudy (ajprincess):

\(36x^2y^6 - 16\) \(=6^2x^2(y^3)^2-4^2\) \(=(6xy^3)^2-4^2\) Difference of squares. \(a^2-b^2=(a-b)(a+b)\) Here \(a=6xy^3\) and \(b=4\) Is it clear?

OpenStudy (anonymous):

Good explanation :) Hope tkennedy is reading it :)

OpenStudy (ajprincess):

thnx a lot for the compliment @JakeV8

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

now can you help me with 9x^4y^2 - z^6

OpenStudy (anonymous):

@tkennedy using the "difference of squares" idea that you just saw eloquently explained, can you make this next problem into a difference of squares format? May need to play with the exponents...

OpenStudy (ajprincess):

\(36x^2y^6-16\) \(=4(9x^2(y^3)^2-4)\) \(=4(3^2x^2(y^3)^2-2^2)\) \(=4((3xy^3)^2-2^2\) Difference of squares \(a^2-b^2=(a-b)(a+b)\) here \(a=3xy^3\) and \(b=2\)

OpenStudy (anonymous):

yes thank you. Now 1 - 100x2

OpenStudy (ajprincess):

I remember helping u to solve this earlier. Am I right?

OpenStudy (anonymous):

i don't remember

OpenStudy (anonymous):

how about a^2b^2 - d^2

OpenStudy (ajprincess):

Can u plz take a look at this http://openstudy.com/users/tkennedy#/updates/5061e8f9e4b0583d5cd2de00

OpenStudy (ajprincess):

What problem do u hav in "difference of squares" questions?